OpenStudy (mikezack123):

A total of 624 tickets were sold for the school play. They were either adult tickets or student tickets. There were 74 more student tickets sold than adult tickets. How many adult tickets were sold?

OpenStudy (moazzam07):

ok i would solve this like this. 624 divided by 2 = 312 and it says 74 more students tickets were sold than the adults then i would do 312 - 74 = 238 so the answer i would say is (There were 238 Adult tickets sold)

OpenStudy (mikezack123):

Thank you so much! :)

OpenStudy (mikezack123):

its not correct :(

OpenStudy (moazzam07):

ok let me see again

OpenStudy (mikezack123):

@thomaster please help? :)

OpenStudy (mikezack123):

@pooja195

OpenStudy (mikezack123):

@mathstudent55

OpenStudy (mikezack123):

@undeadknight26 @jagr2713

OpenStudy (mikezack123):

@Data_LG2 @welshfella @Elsa213

OpenStudy (welshfella):

let x = adult tickets sold number student tickets sold = x + 74 so we have x + x + 74 = 624 solve for x

OpenStudy (mikezack123):

x=275

OpenStudy (welshfella):

yep

OpenStudy (mikezack123):

is that all?

OpenStudy (mikezack123):

Thank you!

OpenStudy (welshfella):

thats it 275 adult tickets sold

OpenStudy (welshfella):

yw

OpenStudy (mikezack123):

can you help me with another please?

OpenStudy (welshfella):

these type of problems are usually best done using algebra

OpenStudy (welshfella):

ok

OpenStudy (mikezack123):

A local hamburger shop sold a combined total of 568 hamburgers and cheeseburgers on Thursday. There were 68 more cheeseburgers sold than hamburgers. How many hamburgers were sold on Thursday?

OpenStudy (mikezack123):

@welshfella

OpenStudy (mikezack123):

nvm i got it, the answer is 250. thanks for your help! the questions were similar :)

OpenStudy (welshfella):

ok yes u have right answer